# Function-Space Decoupled Diffusion for Forward and Inverse Modeling in Carbon Capture and Storage

### v1

@@ 1 Introduction
To bridge this gap, generative diffusion models have emerged as powerful priors for scientific inverse problems (Song et al., 2020; Chung et al., 2022; Yang et al., 2023; Yao et al., 2025). However, applying them to CCS encounters a critical limitation in existing joint-state architectures, which learn the joint distribution of geological parameters and dynamic states $p(\bm{m},\bm{s})$ together. We observe that joint training often leads to physical inconsistency, as the model learns statistical correlations rather than explicit physical laws, especially when paired training data is limited (Lin et al., 2025).
@@ 2 Related Work
Diffusion Models for Inverse Problems. Denoising Diffusion Probabilistic Models (DDPMs) (Ho et al., 2020) currently define the state-of-the-art in generative modeling. For inverse problems, methods such as Diffusion Posterior Sampling (DPS) (Chung et al., 2022; Huang et al., 2024) allow for conditional generation by using the gradient of a forward operator to guide the reverse diffusion process. In the scientific domain, this has been extended to infinite-dimensional function spaces (Lim et al., 2023; Kovachki et al., 2021). Most relevant to our work is the Decoupled Diffusion Inverse Solver (DDIS) (Lin et al., 2025), which showed the failure modes of joint-state training. We extend this decoupled idea to function-space CCS modeling, specifically addressing the extreme sparsity of dynamic monitoring data ($1\%$ coverage) and providing a rigorous quantitative benchmark against the ground truth posterior obtained via Rejection Sampling.
@@ 3.1 Problem Formulation
$p(\bm{m}|\bm{y}_{obs})\propto p(\bm{y}_{obs}|\bm{m})p(\bm{m})$ (2)
Our decoupled approach addresses this by learning a prior $p(\bm{m})$ independently and using a pre-trained neural operator surrogate to efficiently approximate the likelihood $p(\bm{y}_{obs}|\bm{m})$ during posterior sampling, allowing for robust posterior sampling even under extreme data sparsity.
@@ 3.2 Function-Space Diffusion Prior
where $\lambda(\sigma)$ is a noise-level dependent weighting function. Once trained, unconditional sampling from the prior $p(\bm{m})$ is performed via the probability flow ODE:
@@ 3.3 Fun-DDPS: Decoupled Posterior Sampling
To sample from the posterior, we employ Diffusion Posterior Sampling (DPS) (Chung et al., 2022), which guides the reverse diffusion process using the gradient of the log-likelihood. Appendix A provides unified pseudocode and implementation details for our posterior sampler. The conditional score decomposes into the prior score (provided by $\bm{D}_{\theta}$) and the likelihood guidance:
$\nabla_{\bm{m}}\log p(\bm{m}|\bm{y}_{obs})=\overbrace{\nabla_{\bm{m}}\log p(\bm{m})}^{\text{Prior (via }\bm{D}_{\theta})}+\overbrace{\nabla_{\bm{m}}\log p(\bm{y}_{obs}|\bm{m})}^{\text{Likelihood guidance}}$ (6)
The form of the likelihood guidance depends on what is observed.
@@
Eq. 8 represents the core innovation of Fun-DDPS. By separating the prior $p(\bm{m})$ from the physics $\mathcal{L}_{\phi}$, we ensure that the generated samples strictly adhere to the geological manifold by the surrogate. This mechanism is visualized in Figure 2, which highlights how Fun-DDPS translates sparse constraints in the solution space into dense guidance in the parameter space via the surrogate’s Jacobian. This contrasts with joint-state approaches (Fun-DPS, Figure 2b; see Appendix B for details), where gradients from sparse observations remain localized to the specific channels, often resulting in high-frequency artifacts and physical inconsistencies when training data is limited.
@@ 3.4.2 Surrogate Architecture (LocalNO)
For the forward surrogate $\mathcal{L}_{\phi}$, we employ a Local Neural Operator (LNO) (Liu-Schiaffini et al., 2024) to learn the mapping from the geomodel $\bm{m}$ to the dynamic state $\bm{s}$. The training set consists of 12,000 pairs $\{(\bm{m}_{i},\bm{s}_{i})\}$. While standard operators such as FNO (Li et al., 2020) excel at capturing global dependencies, they often produce ringing artifacts around sharp discontinuities. The LNO addresses this by combining global Fourier spectral layers with localized kernels implemented via Discrete Continuous Convolutions (DISCO). This dual-path design is critical for our application: the Fourier path resolves global pressure responses, while the DISCO path accurately captures the shock-like saturation fronts characteristic of multiphase flow. We note that the surrogate is relatively robust to train, especially when paired data is less; and our sampling pipeline is tolerant to surrogate approximation error.
@@ 4 Experiments
We evaluate the proposed Fun-DDPS framework on two distinct tasks: (1) forward modeling conditioned on partial static inputs (geomodel), and (2) inverse modeling conditioned on partial dynamic observations (saturation), as schmatically shown in Fig. 3. For the forward problem, we benchmark performance against two baselines: Fun-DPS: A joint-state diffusion baseline; and Surrogate ($\mathcal{L}_{\phi}$): A deterministic neural operator trained on complete inputs. For the inverse problem, we provide rigorous validation by comparing our approximate posteriors against a “ground truth” posterior distribution obtained via Rejection Sampling (RS) (Robert et al., 1999).
@@ 4.1.1 Problem Setup
In this task, the goal is to predict the dynamic saturation field $\bm{s}$ given a partially observed geomodel $\bm{y}_{geo}$. We evaluate performance across three observation sparsity levels: 100% (full), 50% (random), and 25% (random) coverage.
The primary challenge in this regime is that deterministic surrogates ($\mathcal{L}_{\phi}$) require dense, complete inputs. Standard approaches, such as zero-filling missing values, push the input out of the training distribution, typically leading to severe prediction errors. Fun-DDPS addresses this by utilizing the learned diffusion prior to probabilistically reconstruct missing geomodel features before the forward simulation.
@@ 4.1.2 Results
Qualitative results are visualized in Figure 4. While the surrogate produces noisy, incoherent predictions in the presence of gaps, Fun-DDPS maintains coherent spatial structures, particularly capturing the sharp gradients at the CO2 plume front. Notably, the Fun-DPS baseline shows consistently high error (34–42%) regardless of sparsity. This performance gap highlights a fundamental advantage of the proposed decoupled architecture over joint-state diffusion models. While joint-state models (Fun-DPS) implicitly learns the physical coupling between geomodel and saturation fields as statistical correlations within the joint distribution, they often struggle to capture precise physical laws under limited supervision.
@@ 4.2.1 Problem Setup
We design the experimental setup to reflect the extreme data sparsity encountered in practical Carbon Capture and Storage (CCS) projects. In real-world scenarios, subsurface data is often restricted to well logs, which provide only 1D vertical profiles (column observations) within a vast domain. To mimic this challenge, observations are collected from just two monitoring wells: one at the injector and one 491 meters away (Figure 5). Each well provides 64 dynamic data point, totaling 128 measurements. This configuration corresponds to observing less than 1% of the spatial domain, imposing an extreme challenge for inverse methods to resolve the heterogeneous geomodel from such limited information. We add Gaussian noise with $\sigma_{obs}=0.04$ to the observations.
@@ C.2 Problem Setup
To quantitatively measure the discrepancy between the posterior distributions obtained from CCS-DPS and the RS benchmark, we use the Jensen-Shannon (JS) divergence. The JS divergence is a symmetric and smoothed version of the Kullback-Leibler (KL) divergence Bishop & Nasrabadi (2006), providing a bounded metric for the similarity between two probability distributions. It is defined as:

### v2

@@ 1 Introduction
To bridge this gap, generative diffusion models have emerged as powerful priors for scientific inverse problems (Song et al., 2020; Chung et al., 2022; Yang et al., 2023; Yao et al., 2025). However, applying them to CCS encounters a critical limitation in existing joint-state architectures, which learn the joint distribution of geological parameters and dynamic states $p(\bm{m},\bm{s})$ together. We observe that joint training often leads to physical inconsistency, as the model learns statistical correlations rather than explicit physical laws, especially when paired training data is limited (Lin et al., 2026).
@@ 2 Related Work
Diffusion Models for Inverse Problems. Denoising Diffusion Probabilistic Models (DDPMs) (Ho et al., 2020) currently define the state-of-the-art in generative modeling. For inverse problems, methods such as Diffusion Posterior Sampling (DPS) (Chung et al., 2022; Huang et al., 2024) allow for conditional generation by using the gradient of a forward operator to guide the reverse diffusion process. In the scientific domain, this has been extended to infinite-dimensional function spaces (Lim et al., 2023; Kovachki et al., 2021). Most relevant to our work is the Decoupled Diffusion Inverse Solver (DDIS) (Lin et al., 2026), which showed the failure modes of joint-state training. We extend this decoupled idea to function-space CCS modeling, specifically addressing the extreme sparsity of dynamic monitoring data ($1\%$ coverage) and providing a rigorous quantitative benchmark against the ground truth posterior obtained via Rejection Sampling.
@@ 3.1 Problem Formulation
$p(\bm{m}|\bm{y}_{obs})\propto p(\bm{y}_{obs}|\bm{m})p(\bm{m}),$ (2)
where $p(\bm{m})$ denotes the Radon–Nikodym derivative of the prior measure with respect to a Gaussian reference measure on $\mathcal{M}$. Our decoupled approach addresses this by learning a prior $p(\bm{m})$ independently and using a pre-trained neural operator surrogate to efficiently approximate the likelihood $p(\bm{y}_{obs}|\bm{m})$ during posterior sampling, allowing for robust posterior sampling even under extreme data sparsity.
@@ 3.2 Function-Space Diffusion Prior
where we choose $\|\cdot\|_{\mathcal{M}}$ to be $L_{2}$ norm, and $\lambda(\sigma)$ is a noise-level dependent weighting function. Once trained, unconditional sampling from the prior $p(\bm{m})$ is performed via the probability flow ODE:
@@ 3.3 Fun-DDPS: Decoupled Posterior Sampling
To sample from the posterior, we follow Fun-DPS (Yao et al., 2025), which guides the reverse diffusion process in function space using the gradient of the log-likelihood. Appendix A provides unified pseudocode and implementation details for our posterior sampler. The conditional score decomposes into the prior score (provided by $\bm{D}_{\theta}$) and the likelihood guidance:
$\nabla_{\bm{m}}\log p(\bm{m}|\bm{y}_{obs})=\overbrace{\nabla_{\bm{m}}\log p(\bm{m})}^{\text{Prior (via }\bm{D}_{\theta})}+\overbrace{\nabla_{\bm{m}}\log p(\bm{y}_{obs}|\bm{m})}^{\text{Likelihood guidance}},$ (6)
where $\nabla_{\bm{m}}$ denotes the Fréchet derivative of the corresponding functional. The form of the likelihood guidance depends on what is observed.
@@
Eq. 8 represents the core innovation of Fun-DDPS. By separating the prior $p(\bm{m})$ from the physics $\mathcal{L}_{\phi}$, generated samples remain on the learned geological manifold while data consistency is enforced through surrogate guidance. This mechanism is visualized in Figure 2, which highlights how Fun-DDPS translates sparse constraints in the solution space into dense guidance in the parameter space via the surrogate’s Jacobian. This contrasts with some joint-state approaches (e.g. Fun-DAPS in Lin et al. (2026)), where gradients from sparse observations remain localized to the specific channels, often resulting in high-frequency artifacts and physical inconsistencies when training data is limited.
@@ 3.4.2 Surrogate Architecture (LocalNO)
For the forward surrogate $\mathcal{L}_{\phi}$, we employ a Local Neural Operator (LNO) (Liu-Schiaffini et al., 2024) to learn the mapping from the geomodel $\bm{m}$ to the dynamic state $\bm{s}$. The training set consists of 12,000 pairs $\{(\bm{m}_{i},\bm{s}_{i})\}$. While standard operators such as FNO (Li et al., 2020) excel at capturing global dependencies, they often produce ringing artifacts around sharp discontinuities. The LNO addresses this by combining global Fourier spectral layers with localized kernels implemented via Discrete Continuous Convolutions (DISCO). This dual-path design is critical for our application: the Fourier path resolves global pressure responses, while the DISCO path accurately captures the shock-like saturation fronts characteristic of multiphase flow. We note that the surrogate is relatively robust to train even when paired data are limited, and our sampling pipeline is tolerant to surrogate approximation error.
@@ 4 Experiments
We evaluate the proposed Fun-DDPS framework on two distinct tasks: (1) forward modeling conditioned on partial static inputs (geomodel), and (2) inverse modeling conditioned on partial dynamic observations (saturation), as schematically shown in Fig. 3. In practical deployment, the observation sparsity ratio is usually unknown a priori and can vary across settings. To reflect this, our training pipeline is agnostic to the observation-mask ratio, and sparsity is imposed only at inference time for evaluation. For the forward problem, we benchmark performance against two baselines: Fun-DPS: A joint-state diffusion baseline; and Surrogate ($\mathcal{L}_{\phi}$): A deterministic neural operator trained on complete inputs. For the inverse problem, we provide rigorous validation by comparing our approximate posteriors against a “ground truth” posterior distribution obtained via Rejection Sampling (RS) (Robert et al., 1999).
@@ 4.1.1 Problem Setup
In this task, the goal is to predict the dynamic saturation field $\bm{s}$ given a partially observed geomodel $\bm{y}_{geo}$. We evaluate performance across three observation sparsity levels: 100% (full), 50% (random), and 25% (random) coverage. The forward/inverse task setup is shown in Figure 3.
The primary challenge here is that deterministic surrogates ($\mathcal{L}_{\phi}$) require dense, complete inputs. Standard approaches, such as zero-filling, push the input out of the training distribution, typically leading to severe prediction errors. Fun-DDPS addresses this by utilizing the learned diffusion prior to probabilistically reconstruct missing geomodel features before the forward simulation.
@@ 4.1.2 Results
Qualitative results are visualized in Figure 4. While the surrogate produces noisy, incoherent predictions in the presence of gaps, Fun-DDPS maintains coherent spatial structures, particularly capturing the sharp gradients at the CO2 plume front. Notably, the Fun-DPS baseline shows consistently high error (34–42%) regardless of sparsity. This performance gap highlights a fundamental advantage of the proposed decoupled architecture over joint-state diffusion models. While joint-state models primarily encode geomodel–saturation coupling as statistical correlation in the joint distribution and enforce PDE only at inference time via a physics loss, they can both miss physical laws under limited supervision and have higher sampling cost from physics-loss evaluations.
@@ 4.2.1 Problem Setup
We design the experimental setup to reflect the extreme data sparsity encountered in practical Carbon Capture and Storage (CCS) projects. In real-world scenarios, subsurface data is often restricted to well logs, which provide only 1D vertical profiles (column observations) within a vast domain. To mimic this challenge, observations are collected from just two monitoring wells: one at the injector and one 491 meters away (Figure 5). Each well provides 64 dynamic data points, totaling 128 measurements. This configuration corresponds to observing less than 1% of the spatial domain, imposing an extreme challenge for inverse methods to resolve the heterogeneous geomodel from such limited information. We add Gaussian noise with $\sigma_{obs}=0.04$ to the observations.
@@ 5 Conclusion
Acknowledgments
Anima Anandkumar is supported in part by Bren endowed chair, ONR (MURI grant N00014-23-1-2654), and the AI2050 senior fellow program at Schmidt Sciences. Gege Wen acknowledges the generous support of Schmidt Sciences through the AI2050 fellowship.
@@ C.2 Problem Setup
To quantitatively measure the discrepancy between the posterior distributions and the RS ground truth, we use the Jensen-Shannon (JS) divergence. The JS divergence is a symmetric and smoothed version of the Kullback-Leibler (KL) divergence Bishop & Nasrabadi (2006), providing a bounded metric for the similarity between two probability distributions. It is defined as: